Two coins are tossed. Let T be the number of tails that occur. Determine

the values of the random varable T


1
Expert's answer
2021-04-15T16:55:18-0400

When we toss two coins, the sample space will be:

S = (Head, Head), (Head, Tail), (Tail, Head), (Tail, Tail)

Suppose we let T, be the number of tails that occur;

T=0 when no tails occur

P(T=0)=1/4

T=1 When only one tail occurs. That is, when the outcome is (Head, Tail) or (Tail, Head)

P(T=1)=1/2

T=2, when only two tails occur. When the outcome is (Tail, Tail)

P(T=2)=1/4

Therefore, the random variable T takes the values 0,1, 2